Sam is 12 years younger than Rocco. How old will Rocco be 10 years from now if Sam is now 4 years old?

Answer by stanbon(75887)

Sam is 12 years younger than Rocco. How old will Rocco be 10 years from now if Sam is now 4 years old?
-----
s = r-12
s = 4
----
Now::
s = 4 yrs
r = s+12 = 16 yrs
-----
10 years from now
Rocco will be 16+10 = 26 yrs
---------------------------
Cheers,
Stan H.
----------